OpenStudy (anonymous):

There is a circular pool in the backyard with an approximate area of 113 square inches. What was the diameter?

OpenStudy (anonymous):

\[113= \pi r ^{2}= \pi (d/2)^{2}\] \[113= \frac{ \pi d^{2} }{ 4 }\] you can solve for d now.

OpenStudy (shamim):

Area of a circle is\[A=\pi \times r ^{2}= \pi \times (\frac{ d }{ 2 })^{2}\]

OpenStudy (anonymous):

still dont understand 1111111111:{

OpenStudy (shamim):

ok

OpenStudy (anonymous):

here, r is the radius d is the diameter. we know that, radius is the half of the diameter of the circle. r=d/2

OpenStudy (shamim):

\[113= \pi \times \frac{ d ^{2} }{ 4 }\]

OpenStudy (shamim):

\[d ^{2}=\frac{ 113 \times 4 }{ \pi }\]

OpenStudy (anonymous):

im still geting 28.25 -.-

OpenStudy (anonymous):

ooooooooh